### v1

Abstract
Replacing fossil-fueled appliances and vehicles with electric alternatives can reduce greenhouse gas emissions and air pollution in many settings. However, residential electrification can raise electricity demand beyond the safe limits of electrical infrastructure, increasing the risk of blackouts or requiring grid reinforcement that can be slow and expensive. Here, we estimate the physical and economic impacts on distribution grids of electrifying all housing and personal vehicles in each county of the lower 48 United States. We find that space heating is the main driver of grid impacts, with the coldest regions seeing demand peaks up to three times higher than today’s peaks. Accommodating electrification of all housing and personal vehicles could require up to 312 GW of distribution grid reinforcement nationally, at a cost of \$183 to \$415 billion, or \$1,500 to \$3,400 per household (95% confidence intervals). However, demand-side management can mitigate demand peaks, reducing grid reinforcement costs by up to 92%.
keywords:
Distribution grid, Electrification, Heat pumps, Electric vehicles
Fossil-fueled driving, space heating, and water heating contribute about one-third of United States greenhouse gas emissions [1]. Replacing fossil-fueled vehicles and heaters with electric alternatives could significantly reduce these emissions today [2], and essentially eliminate them if the United States achieves its goal of 100% carbon-free electricity by 2035 [3]. For these reasons, the Inflation Reduction Act of 2022 incentivizes adoption of electric vehicles and heat pumps (efficient electric heating and cooling machines) through tax credits and point-of-sale rebates [4, 5, 6]. State and local governments have enacted further incentives, including some bans on fossil-fueled vehicles or heaters [7]. Due in part to these policies, adoption of electric vehicles and heat pumps in the United States is accelerating [8].
While electrification of fossil-fueled activities could deeply reduce emissions, it could also have serious impacts on the power grid. Today, electricity demand in most United States power systems peaks in the hottest hours of the year, driven mainly by air conditioning [9]. Electrifying space heating could heighten annual demand peaks and shift them from summer to winter [10, 11, 12, 13, 14]. Electric vehicle adoption could further exacerbate demand peaks [15, 16, 17]. Heightened demand peaks from electrification could require replacing power lines and transformers, which must be sized to accommodate worst-case power flows. Transformers in particular are already in short supply, with United States utilities experiencing lead times of up to two years and prices that have increased by 400 to 900% in the last three years [18]. Utilities pass grid reinforcement costs on to ratepayers in the form of increased electricity prices [19, 20], which could weaken economic incentives for further electrification.
Existing research on nationwide grid impacts of electrification focuses mainly on high-voltage transmission systems. The National Renewable Energy Laboratory’s Electrification Futures Study [21, 10, 22] investigates how electrification might alter load profiles and transmission infrastructure needs. In their high-electrification 2050 scenario, peak electricity demand rises 38% relative to a reference scenario. Wilson et al. [2] use the ResStock tool [23] to investigate cost and emission impacts under various heat pump adoption scenarios. While [2] does not explicitly consider grid impacts, Wilson et al. stress the need to understand how evolving load profiles should shape grid infrastructure investments.
In parallel with nationwide transmission-level research, several studies investigate distribution-level grid impacts in California. Notably, Elmallah et al. [24], use the Distribution Deferral Opportunity Report, which the California Public Utilities Commission requires utilities to file each year, to analyze electrification impacts within the Pacific Gas & Electric service territory. Elmallah et al. estimate that distribution grid reinforcement to accommodate new load from electric vehicles and heat pumps could cost \$1 to \$10 billion, or about \$200 to \$2,000 per customer. Li et al. [25] use the same dataset to study electric vehicle adoption statewide, estimating that 67% of distribution feeders will need reinforcement, at a total cost of \$6 to \$20 billion. Several other state-specific studies have investigated the impacts of electrification on emissions [14, 26, 27].
As a complement to nationwide transmission-level studies and California-specific distribution-level studies, this paper investigates the impacts of electrification of space heating, water heating, and personal vehicles on distribution grids nationwide. We use the open-source EDGIE (Emulating the Distribution Grid Impacts of Electrification) toolbox [28] and make all data publicly available to facilitate reproduction or extension of the work. As illustrated in Fig. 1, EDGIE models heat pumps, electric vehicles, water heaters, and building thermal dynamics. It uses real data on weather [29], miscellaneous electrical loads [30], indoor temperature setpoints [31], and domestic hot water use [32]. EDGIE implements empirically validated, multi-physics modeling at hourly time resolution and county-level spatial resolution. We tune building, appliance, and vehicle properties to county-specific data. We find that almost all United States counties could require grid reinforcement in a 100% electrification scenario. In colder climates, peak demand rises by 180% on average relative to today’s peak demand, and up to 300% in the coldest counties. Distribution grid reinforcement to accommodate 100% electrification nationwide could cost \$183 to \$415 billion. However, demand-side management strategies, such as improving equipment efficiencies and coordinating device operation, could reduce nationwide grid reinforcement costs by up to 92%.
Figure 1: The open-source EDGIE toolbox uses high-resolution, bottom-up modeling to estimate the impacts of home and vehicle electrification on distribution grids in each county of the lower 48 United States.
Impacts on physical infrastructure
We assess the physical impacts of home and vehicle electrification on distribution grids through bottom-up modeling. In each county, we simulate all the residential appliances and personal vehicles from 1,000 representative households, then rescale the aggregate load profile to reflect the true number of households. Simulated appliance and vehicle properties, housing sizes and types (detached single-family vs. attached multifamily), and levels of insulation and air sealing all vary both across and within counties. We simulate peak heating and cooling weeks under two scenarios: (1) A business-as-usual scenario that reflects today’s building stock, appliances, and vehicles, and (2) an all-electric scenario that reflects a possible future with complete electrification of personal vehicles and residential space and water heating. In the all-electric scenario, each home has a resistance water heater, an air-to-air heat pump for space heating and cooling, and backup resistance heat. We compute design heating and cooling loads for each home under EnergyStar heating and cooling design conditions [33]. Each heat pump is sized to meet the larger of its design heating load and design cooling load, capped at the largest residential air-to-air heat pump size currently available. In the two northernmost climate zones, we simulate adoption of heat pumps that meet the United States Department of Energy’s Cold Climate Heat Pump Challenge specifications [34]. We fit physical and behavioral parameters associated with buildings, thermal equipment, and vehicles to county- and climate-specific data, using real data wherever possible. Section Methods and the Supplementary Information document provide more details on models and data sources.
Figure 2: Peak aggregate electricity demand from 1,000 households in a hot week in Miami (left) barely exceeds today’s peak. In a cold week in Minneapolis (right), however, demand peaks at more than double today’s peak.
The modeling approach outlined above enables analysis of the sizes and causes of likely grid impacts from electrification. Fig. 2 shows the aggregate power used by 1,000 households over a peak cooling week in Miami, Florida (left plot; a hot, humid climate) and a peak heating week in Minneapolis, Minnesota (right; cold). In these plots, the shaded areas from bottom to top represent miscellaneous electrical loads (blue; lights, computers, dishwashers, etc.), water heaters (orange), electric vehicles (yellow), and heat pumps with backup resistance heat (purple). Dashed red lines show today’s annual demand peaks. In Miami, electrification increases peak demand only modestly. Electric vehicle adoption increases electricity demand, but the new air-to-air heat pumps – which can both heat and cool – offset this increase substantially by improving air conditioning efficiency. In Minneapolis, by contrast, peak demand more than doubles due to space heating.
Figure 3: Today (top left), per-household demand peaks are typically higher in hotter areas. In an all-electric future (top right), per-household peaks increase and shift to colder areas. Per-household grid reinforcement requirements (bottom) are highest in the Rocky Mountains, northern Midwest, and Northeast.
We analyze grid impacts nationwide by running simulations for each county in the lower 48 United States. Fig. 3 shows heat maps of the estimated distribution grid capacity, normalized by the number of homes, in each county under the business-as-usual scenario (top left) and the all-electric scenario (top right). We estimate distribution grid capacities in each county and scenario by simulating the appliances and vehicles for 1,000 households, rescaling the aggregate demand to the true number of households, taking the 99th percentile, then adding 20% to reflect an infrastructure sizing safety margin. The top left plot shows that the normalized grid capacity (in units of kW per household) in the business-as-usual scenario tends to be higher in areas with more cooling demand, reaching a maximum of 13 kW per household in the hot, sunny Southwest. By contrast, the top right plot shows that the normalized grid capacity in the all-electric scenario is higher in colder areas, reaching a maximum of 25 kW per household in the northern Midwest.
The lower heat map in Fig. 3 shows the estimated grid reinforcement capacity, normalized by the number of homes, required to accommodate 100% electrification in each county. We estimate the grid reinforcement capacity requirement by the difference between the all-electric and business-as-usual grid capacities. As expected, northern regions require more grid reinforcement due to higher space heating demand. Grid reinforcement requirements reach a maximum of 21 kW per household in the northern Midwest. Many southern areas with mild winters require little or no grid reinforcement. In some cases, the lower heat map in Fig. 3 shows significant variation in grid reinforcement requirements between adjacent counties. This variation can be explained by differences in housing types and sizes, levels of insulation and air sealing, driving patterns, weather extremes, or other location-specific data. These geographic differences suggest a need for tailored infrastructure planning, as approaches that work in one county may not necessarily work in neighboring counties.
Distribution grid reinforcement costs
To assess the economic costs of reinforcing distribution grids to accommodate electrification, we multiply the aggregate grid reinforcement requirement for each county (in units of kW) by a \$/kW price. Informed by several recent studies, we choose this price to reflect a typical distribution grid upgrade. In reality, grid reinforcement prices vary significantly across the United States due to variation in grid topology and capacity, prevalence of overhead vs. underground power lines, costs of labor and equipment, and other factors. Due to a lack of reliable location-specific data on grid reinforcement prices, however, we use the same price distribution for each county.
We draw on four studies [24, 35, 36, 37] to determine a plausible range of prices for typical distribution grid reinforcement, which may include installing power lines, transformers, switchgear, capacity banks, or other infrastructure. Based on Pacific Gas & Electric data, Elmallah et al. [24] estimate prices of 860 to 36,500 \$/kW, with the highest prices for the smallest projects [^1] We inflation-adjusted all price data to 2024 dollars.. Larson et al. [35] use Energy Information Administration data [38] to derive state-wise prices and a national average of 1,664 \$/kW, including a variety of administrative costs. Fares et al. [36] estimate an average price of 69 \$/kW based on data from the Federal Energy Regulatory Commission’s Form 1, where utilities can report their capital, operation, and maintenance costs related to grid reinforcement. In estimating this price, Fares et al. delineate distribution grid reinforcements from generation, transmission, and administrative costs. Rauschkolb et al. [37] use the same Federal Energy Regulatory Commission dataset to develop a linear model that accounts for both one-time capital costs of expanding grid capacity, and for increases in the annual costs of operating and maintaining a larger grid.
Here, we modify the approach of Rauschkolb et al. to estimate a representative distribution grid reinforcement price. We assume that grid capacities increase linearly from the business-as-usual capacities in 2025 to the all-electric capacities in 2050. We map future costs into net present costs using discounted cash flow analysis. As in [37], the modified approach accounts for both one-time capital costs of reinforcing infrastructure, and increased annual costs from operating and maintaining a larger grid. The Supplementary Information document contains the details of our modifications, which yield a mean grid reinforcement price of 960 \$/kW and a 95% confidence interval of 587 to 1,331 \$/kW.
The upper row of heat maps in Fig. 4 shows the estimated total distribution grid reinforcement cost for each of the lower 48 United States under low (left), medium (middle), and high (right) price scenarios. We compute these estimates by multiplying the grid reinforcement requirement, aggregated over all counties in the state, by prices of 587, 960, and 1,331 \$/kW, corresponding to the 2.5th, 50th, and 97.5th price percentiles. Unlike the per-household grid reinforcement requirement estimates in Fig. 3, the aggregate cost estimates in Fig. 4 reflect population density as well as load growth intensity. States with both high populations and high per-household grid reinforcement requirements have the highest total costs. For example, the 7.6 million households in New York, a densely populated state with cold winters, see a total cost of about \$30 billion, or about \$4,000 per household. Per-household grid reinforcement costs, shown in the bottom row of heat maps in Fig. 4, reach a maximum of \$16,000 per household under the 97.5th price percentile in the cold, sparsely populated state of North Dakota.
Aggregated over the lower 48 United States, grid reinforcement to accommodate electrification could cost \$183 to \$415 billion, or \$1,500 to \$3,400 per household. For context, United States households spend about \$233 billion (\$1,900 per household) per year on residential energy bills [39]. The Inflation Reduction Act of 2022 allocates about \$370 billion over ten years to spending on energy and climate [40]. The United States annual military budget was about \$816 billion in 2023 [41].
Figure 4: Statewide distribution grid reinforcement costs (top row) reflect both population densities and per-household grid reinforcement requirements. Per-household costs (bottom row) do not reflect population densities. Left column: Low-price scenario for grid reinforcement. Middle: Average price. Right: High price.
Cost reductions from smart electrification
The estimated distribution grid reinforcement requirements and costs in the previous sections come from an electrification scenario that extrapolates forward today’s housing stock, vehicle fleet, equipment configurations and efficiencies, and user behavior. Unmodified, these electrification choices could require \$183 to \$415 billion in distribution grid reinforcement costs. However, electrification with strategic demand-side management – referred to here as ‘smart electrification’ – could reduce these costs substantially.
To assess the cost reduction potential of smart electrification, we simulate four demand-side management strategies. First, the resistance water heaters simulated above could shift to heat-pump water heaters, which use about half the electricity per unit of heat output. Second, thermal demand from buildings could decrease due to better insulation and air sealing, due to shifts from detached single-family housing to less energy-intensive attached multifamily housing, or due to shifts to smaller housing. Third, software could reduce peak aggregate demand by coordinating the operation of water heaters, electric vehicles, and heat pumps. For example, heat pumps could preheat homes in anticipation of aggregate demand peaks, or electric vehicles could interleave their charging with heat pump operation. We simulate device coordination through convex optimization with perfect forecasts in one representative location per state, as discussed in the Supplementary Information document. Fourth, the air-source heat pumps simulated above could shift to ground-source heat pumps, which use about one-third less electricity per unit of heat output and require less backup resistance heat.
Figure 5: Improving equipment efficiencies (top left, bottom right), reducing thermal demand (top right), or coordinating device operation (bottom left) can reduce the grid impacts of electrification.
Fig. 5 shows heat maps of the estimated reductions in grid reinforcement requirements in each county under the four demand-side management strategies. While heat-pump water heaters use much less electricity than resistance water heaters, the top left heat map in Fig. 5 shows that they have little impact on grid reinforcement requirements due to the relatively small contribution of water heating to peak demand. By contrast, the heat maps with less thermal demand from buildings (top right) and with ground-source heat pumps (bottom right) show significant reductions in grid reinforcement requirements, particularly in the coldest regions. The heat map with device coordination (bottom left) shows modest reductions in grid reinforcement requirements. However, these reductions are more evenly distributed across the country because, for example, coordinating electric vehicle charging can reduce peak demand anywhere, regardless of climate.
Figure 6: Grid reinforcement cost reductions from individual (top set of bars) and combined (bottom) demand-side management strategies.
Fig. 6 shows nationwide grid reinforcement cost reductions from the four individual strategies and from combinations thereof. These estimates use the mean grid reinforcement price of 960 \$/kW, which gives a nationwide cost of \$300 billion without demand-side management. Individually (top set of bars), the four strategies reduce costs by 6%, 22%, 52%, and 55%, respectively. Perhaps surprisingly, device coordination – a strategy based mainly on software, sensing, and communication – achieves similar cost reductions to ground-source heat pumps, a significantly more hardware-intensive and expensive strategy. Cost savings from combined strategies (bottom set of bars) are not strictly additive. For example, the combination of heat-pump water heaters (6% individually), less thermal demand from buildings (22%), and ground-source heat pumps (55%) reduces costs by 68%. The combination of all four demand-side management strategies reduces nationwide grid reinforcement costs by 92%, to \$25 billion.
Discussion
This paper used bottom-up modeling to estimate the physical and economic impacts of electrification of all residential appliances and personal vehicles on distribution grids in each county of the lower 48 United States. The modeling tools are open-source; the cleaned input data and simulation results are free and public. This paper found that distribution grid reinforcement requirements depend mainly on space heating demand, which can triple peak aggregate demand in the coldest counties relative to business as usual. Without demand-side management, nationwide grid reinforcements could cost \$183 to \$415 billion, or \$1,500 to \$3,400 per household, but smart electrification could reduce these costs by up to 92%.
This paper’s results could find use by grid planners, for whom spatially-resolved estimates of grid reinforcement requirements can inform capacity expansion decisions. The demand-side management results could also find use by policymakers who determine incentives and regulations that shape individuals’ electrification choices. For example, the effectiveness of device coordination in reducing grid reinforcement costs could motivate policymakers to encourage or require electrification with appliances and vehicles that have the sensing, communication, and computing capabilities that device coordination systems require.
There are many opportunities to extend this work. This paper’s scope covers distribution grid reinforcement driven by electrification of the residential sector and personal vehicles. Future research could extend this work to include transmission grid reinforcement, generation capacity expansion, and electrification of the commercial and industrial sectors. In particular, estimating transmission and generation costs from residential electrification would be straightforward and useful extensions. Future research could also extend the open-source EDGIE toolbox to incorporate solar photovoltaics, stationary batteries, or thermal energy storage. Although we could not find reliable, spatially-resolved data on distribution grid topologies or reinforcement prices, such data would enable analysis of power quality issues and more precise estimation of distribution grid reinforcement costs. This paper used historical weather data; future work could incorporate the effects of climate change on demand peaks driven by extreme weather. Finally, while this paper estimated the value of demand-side management strategies in mitigating distribution grid reinforcement costs, these strategies could also reduce generation and transmission expansion costs, greenhouse gas emissions, air pollution, and household energy bills. Identifying appropriate, location-specific mixes of grid upgrades and demand-side management measures is an interesting direction for future work.
Methods
[…]
Data
We tune model parameters to match real data wherever possible. Distributions of building and appliance properties follow the ResStock database [42]. Heat pump coefficient of performance curves come from manufacturer data aggregated by the Northeast Energy Efficiency Partnerships [43]. We generate domestic hot water time series from [32]. The number of vehicles per household in each county comes from census data [44]. We classify personal vehicles as either small (e.g., coupes and sedans) or large (e.g., pickup trucks and sport utility vehicles) and tune the distribution of small and large vehicles in each county to survey data from [45]. We tune the driving efficiencies of small and large electric vehicles to the Tesla Model S and Ford F-150 lightning, respectively. For both vehicle types, driving efficiencies vary with outdoor temperatures according to [46]; this variation reflects energy use for cabin heating and cooling, as well as temperature-dependent battery chemistry. Commute distances follow census data for each county [47]. The Supplementary Information document contains further details on modeling, simulation, and data sources. Table 1 summarizes the data sources drawn upon by this paper.
Table 1: Data Sources
Parameter Data Source References
Building types and floor areas National Renewable Energy Laboratory’s ResStock [23, 48]
Business-as-usual appliances National Renewable Energy Laboratory’s ResStock [23, 48]
Building insulation and air sealing ASHRAE code [49]
Design heating and cooling temperatures EnergyStar [33]
Heat pump coefficients of performance Northeast Energy Efficiency Partnerships [43]
Domestic hot water profiles Building America analysis spreadsheets [32]
Commute times Census survey [50]
Vehicles per household Census survey [50]
Electric vehicle parameters Manufacturer specification and journals [46, 51]
Weather data Oikolab [29]
Climate zones ASHRAE/IECC [52]
Grid upgrade prices Journals [37]
[…]
Space heating and cooling EDGIE models indoor temperature dynamics using first-order linear ordinary differential equations:
$C_{1}\dot{T}_{1}(t)=\frac{\theta_{1}(t)-T_{1}(t)}{R_{1}}+q_{1}(t)+w_{1}(t)$ (1)
$0\leq q_{1}(t)\leq\eta_{1}(\theta_{1}(t))\overline{p}_{1h}+\overline{p}_{1r}.$
Here $t$ (h) denotes time, $T_{1}$ (C) is the indoor air temperature, $C_{1}$ (kWh/C) is the indoor air’s thermal capacitance, $\theta_{1}$ (C) is the outdoor air temperature, $R_{1}$ (C/kW) is the thermal resistance between indoor and outdoor air, $q_{1}$ (kW) is the thermal power supplied by the heat pump and/or resistance heater, $w_{1}$ (kW) is the exogenous thermal power from the sun, plug loads, lights, body heat, etc., $\eta_{1}$ is the heat pump’s coefficient of performance curve, $\overline{p}_{1h}$ (kW) is the heat pump’s electric power capacity, and $\overline{p}_{1r}$ (kW) is the electric power capacity of the resistance heater. In discrete time,
$T_{1}(k+1)$ $=a_{1}T_{1}(k)+(1-a_{1})[\theta_{1}(k)+R_{1}(q_{1}(k)+w_{1}(k))],$ (2)
where $k$ indexes time steps, $a_{1}=\exp(-\Delta t/(R_{1}C_{1}))$, and $\Delta t$ (h) is the time step duration. Backup resistance heat runs only when the heat pump cannot meet heating demand. The heat pump’s control system tries to perfectly track its setpoint $\hat{T}$ (C) by delivering thermal power
$\hat{q}_{1}(k)=\frac{1}{R_{1}}\left(\frac{\hat{T}_{1}(k+1)-a_{1}T_{1}(k)}{1-a_{1}}-\theta_{1}(k)\right)-w_{1}(k),$ (3)
but may saturate at an upper or lower capacity limit:
$p_{1}(k)=\begin{cases}0&\hat{q}_{1}(k)\leq 0\\ \hat{q}_{1}(k)/\eta_{1}(k)&0<\hat{q}_{1}(k)\leq\eta_{1}(k)\overline{p}_{1h}\\ (1-\eta_{1}(k))\overline{p}_{1h}+\hat{q}_{1}(k)&\eta_{1}(k)\overline{p}_{1h}<\hat{q}_{1}(k)\leq\overline{p}_{1h}+\overline{p}_{1r}\\ \overline{p}_{1h}+\overline{p}_{1r}&\overline{p}_{1h}+\overline{p}_{1r}<\hat{q}_{1}(k).\\ \end{cases}$ (4)
Experimental validation of the space heating model can be found in [28]. Space cooling works similarly but without backup resistance heat.
[…]
Electric vehicles EDGIE models electric vehicle batteries via
$\dot{E}(t)=-rE(t)+\eta_{2}p_{2}(t)-w_{2}(t)$ (5)
$0\leq E(t)\leq\overline{E}$
$0\leq p_{2}(t)\leq\overline{p}_{2}.$
Here $E$ (kWh) is the chemical energy stored in the battery, $r$ (1/h) is the battery’s self-dissipation rate, $\eta_{1}$ is the charging efficiency, $p_{1}$ (kW) is the electric charging power, $w_{1}$ (kW) is the chemical power discharged to drive the vehicle, $\overline{E}$ (kWh) is the energy capacity, and $\overline{p}_{1}$ (kW) is the charging power capacity. In discrete time,
$E(k+1)=a_{2}E(k)+\frac{1-a_{2}}{r}(\eta_{2}p_{2}(k)-w_{2}(k)),$ (6)
where $a_{2}=\exp(-r\Delta t)$.
[…]
Water heating EDGIE models water heaters via
$C_{3}\dot{T}_{3}(t)=\frac{\theta_{3}-T_{3}(t)}{R_{3}}+q_{3}(t)-w_{3}(t)$ (7)
$0\leq q_{3}(t)\leq\eta_{3}\overline{p}_{3h}+\overline{p}_{3r}.$
Here $T_{3}$ (C) is the water temperature, $C_{3}$ (kWh/C) is the water’s thermal capacitance, $\theta_{3}$ (C) is the (constant) air temperature surrounding the tank, $R_{3}$ (C/kW) is the thermal resistance between the water and surrounding air, $q_{3}$ (kW) is the thermal power supplied to the tank, $w_{3}$ (kW) is the thermal power withdrawn for showers, dish-washing, laundry, etc., $\eta_{3}$ is the water heater’s coefficient of performance ($\eta_{3}=1$ for resistance water heaters), $\overline{p}_{3h}$ (kW) is the heat pump’s electric power capacity ($\overline{p}_{3h}=0$ for resistance water heaters), and $\overline{p}_{3r}$ (kW) is the electric power capacity of the resistance heater. In discrete time,
$T_{3}(k+1)=a_{3}T_{3}(k)+(1-a_{3})[\theta_{3}+R_{3}(q_{3}(k)-w_{3}(k))],$ (8)
where $a_{3}=\exp(-\Delta t/(R_{3}C_{3}))$. The water heating control logic, including dispatch of backup resistance heat for hybrid heat-pump water heaters, is the same as the space heating control logic.
Declarations
[…]
Data and code availability The data and code that support this study are available at https://github.com/priyada7/EDGIE.
[…]
Supplementary information The online version contains supplementary material available at LINK.